Please help. The length of each side of a cube is increased by 20%, what is the % increase in the surface area of the cube?
Answer by MissionPossible(18)

suppose the length of the small(old) cube is X.
The length of each side of a cube is increased by 20% ,then the length of the bigger(new) cube would be 1.2X (120% = 120*x/100 = 1.2X
The surface of the small(old) cube is 6X^2
The surface of the bigger(new) cube is 6(1.2X)^2 = 1.44* 6X^2
Then the % increase in the surface area of the cube = 44 % Ans